# Words and Action: Modeling Linguistic Leadership in #BlackLivesMatter Communities

### v1

@@ Results
Table 2: Semantic changes Examples of semantic changes with the tweets in which they appear, paraphrased from the original.
@@
Our approach offers insight into the changing meaning of conceptual keywords (?) related to political activism and the BLM movement, as well as about the specific communities that helped to spread those changed meanings. Here we discuss several sets of semantic changes identified by our model and the significance of the overall network we construct. Taken together, these results affirm the validity of our model and its underlying community clustering approach, as well as the crucial work done by BLM activists to shape conversations about police violence, antiblack racism, and possible solutions to both. They also affirm the importance of Black celebrities in communicating the ideas of BLM to the general public, as well as the presence of conservative critics who push back against activist discourse–and, at times, co-opt its language. Below we discuss the findings from each phase of our methodological pipeline, building towards a conclusion about the decade-long lead-up to the current political climate in which the words of antiracism themselves have become the subject of legislative attention.
@@
The first part of our model identifies the words that demonstrated statistically-significant semantic changes (Table 2). Beginning with general terms related to activism, we find that the term activists undergoes a shift in meaning between 2014-2015, following the murders of Freddie Gray and Sandra Bland, when it has no particular connection to the work of BLM–its near neighbors are terms like “detained,” “targeted,” and “scuffle”–to the summer of 2020, by which point the term is very clearly associated with BLM activists in particular, as indicated by near neighbors including “blackled,” “movements,” “bipoc,” and “advocates.” The term action, similarly, enters the dataset with no specific connotations of activism (its near neighbors are related to TV news channels) but, by the summer of 2020, has acquired specific valences of political change, as evidenced by the near neighbors of “meaningful,” “tangible,” “steps,” and “stand.” In a similar manner, the word active also acquires connotations specific to allyship, as indicated by near neighbors of “learning,” “listening,” and “unlearning,” among others. (The terms allyship and learning undergo similar changes, acquiring each other as near neighbors along similar timeframes).
Terms like “listening” and “learning” point to more specific ways that BLM activists shaped the broader conversation about how to engage in activism around racial justice as well as how to contribute as an ally. But additional terms index specific issues and debates. For example, the term abolish begins with associations with the death penalty and, in 2020, acquires an expanded meaning that encompasses the abolition of the police–a key component of the BLM platform. The term abolishing, similarly, acquires the near neighbors of “defunding,” “reallocating,” “reforming,” and “demilitarization” as the summer of 2020 approaches. We also find theory, as in “critical race theory,” coalescing as a keyword during this time, acquiring associations with both “crt” and “ideology.” Taken together, these word changes point to a clear coalescing around the discourse of racial justice that we have come to associate with the movement.
Figure 5: Linguistic leadership The linguistic leadership of key communities including (a) activists, (b) progressives, (c) Black celebrities, and (d) conservatives with their contributions highlighted. The height of each bar is proportional to the overall number of words a community leads (left) or follows (right). The height of the connecting band is proportional to the number of words the community on the left leads the one on right.
@@
Unexpected insights begin to emerge when looking at the leadership network overall. To be sure, we observe the influence of the BLM activists (Figure  5(a)), who claim an outsized number of word changes in comparison to most other communities. The progressive community, which sits slightly closer to the center than the activist community, can count a significant number of word changes as well (Figure  5(b)), and it would make sense that the pattern of the progressives would track that of the BLM activists but less strongly. What is perhaps unexpected is that the conservative community-moreso than the center/left news media-is the most substantial follower of both the activist and progressive communities. While we might expect a more gradual path from the left to the center, and then to the right, our method shows how the conservative community engaged directly with the activist and progressive communities from the start.
@@
Another unexpected but not surprising result was the strong role of the community of Black celebrities in communicating the ideas of the BLM movement to the general public (Figure 5(c)). Indeed, Black celebrities claim 24.73% of all of word changes, and are followed by activists, then center/left news media, and conservatives in high proportion. This finding suggests that the direct association with the BLM movement may mark its activists for pushback-and, at times, outright aggression-by the conservative community, the celebrity status of those in the Black celebrity community may protect them from some of the conservative vitriol, and as a result, allow them to communicate the ideas of the activists to a broader public. We see this in their leadership of terms like activism, which transforms from the specific context of social movements to a word that itself indexes a role of social importance, as indicated by the near neighbors of “importance,” “career,” “societal,” and “impact.” We also see this in the terms that one would seem to associate with the BLM movement itself, such as listen. The connotations of learning and unlearning, as well as amplifying the voices of those with direct experience-as indicated by the near neighbors of “learn,” “unlearn,” “amplify,” and “explain,” are communicated to the center/left news media commmunity from the Black celebrities. In addition, we find this community influencing the community of progressives, passing along the changed meaning of theory, discussed above, as well as the more basic concept of equity. Interestingly, the conservative community at times learns directly from the Black celebrity community. In particular, the term nonracist is introduced by this community, and is adopted directly from the conservative community from there.
@@
A final observation has to do with the role of the conservative community in introducing new word meanings-and not just adopting them (Figure 5(d)). While few of these words are conceptual keywords, their overall number is substantial. This finding confirms the consistent and persistent engagement of these ideas by the conservative community-not only in the wake of the summer of 2020, but from the movement’s start. That these words are not conceptual keywords at once confirms prior findings about how conservatives-as well as white people as a group-demonstrate a lead-lag relationship with BLM activists (?) and shows how they engaged at the level of discourse: through the wide array of words that make up a language.
@@
As far as the prominence of the conservative community in our study, we see significant implications for our understanding of the emergence of the “anti-woke” and “anti-CRT” agenda that has entered legislation in nine US states and counting (?). Whereas the common narrative is that this movement emerged in the aftermath of the summer of 2020, and has since been fueled by the Republican party, our findings show that there has always been direct conservative engagement with the discourse of Black Lives Matter. Moving forward, this finding may reshape the narrative of “conservative backlash” and, instead, point to a story of continual pushback and even outright aggression that met the movement from its start.
Finally, we would like to highlight a key absence in this project, which is the missing data that might otherwise document the crucial role of the young Black activists who fueled the movement from its start. These regular young people were essential in the movement’s early phases, as they were in organizing the on-the-ground protests that took place in every city where a police killing took place. Qualitative researchers have documented their crucial role (??), as have news reports (?), yet their status as regular young people has meant that they lack the protections granted to celebrities or even to adults. As activists, they have been subjected to harassment and intimidation, and in some cases they have even been killed by the police themselves (??). As early as mid-2015, when Freelon et al released their initial findings on the network structure of #BLM, they found that the young Black activists who had been present in their data in 2014 had disappeared eighteen months later (?). Using the usernames of their young Black activist community as seeds, we found similarly that these users had largely disappeared-save for the few who themselves became famous, and joined the Black celebrity community instead.
In her influential theorization of missing data sets, the artist and educator Mimi Onuoha explains how the gaps in datasets provide “cultural and colloquial hints of what is deemed important,” and further observes how “this lack of data typically correlates with issues affecting those who are most vulnerable in that context” (?). In our #BlackLivesMatter dataset and the findings that it has enabled, these young Black activists constitute this precise form of missing data. Thus as we celebrate our findings for how they affirm the role of the Black Lives Matter movement in shaping a larger public conversation, we must also hold space for what our study cannot confirm, which is the equally crucial role played by the activists whose words are missing from our dataset. Even if their words are unrecorded, they are not unremembered.
@@ Conclusion
In this project we describe a method of modeling semantic leadership across a set of communities associated with the #BlackLivesMatter movement, which has been informed by qualitative research on the structure of social media and of Black Twitter in particular. We describe our bespoke approaches to time-binning, community clustering, and connecting communities over time, as well as our adaptation of state-of-the-art approaches to semantic change detection and semantic leadership induction. We find substantial evidence of the leadership role of BLM activists and progressives, as well as Black celebrities. We also find evidence of the sustained engagement of the conservative community with this discourse, suggesting an alternative explanation for how we arrived at the present moment, in which ”anti-woke” and “anti-CRT” bills are being enacted nation-wide. Finally, we recognize the role of the young Black activists who are missing from our data, and hold space for their contributions even as they cannot be identified through quantitative means.
@@
[Uncaptioned image] Figure 6: Top Community Members The most central users of each subcommunity cluster, with color indicating final community grouping. Users with fewer than 3000 followers or with no longer active X accounts are not shown. Note that ranking by in degree leads to certain users (e.g. @realdonaldtrump) being included in communities in which they engage adversarially.
Maxima Additional Candidate Maxima Sep 02, 2014 n/a Dec 25, 2014 Dec 18, 2014 Aug 31, 2015 Mar 14, 2015; Apr 10, 2015; May 04, 2015 Jul 08, 2016 n/a May 27, 2020 n/a Aug 07, 2020 Sep 05, 2020; Sep 25, 2020
Table 3: Maxima Additional candidate maxima that were merged into the final time bin segments.
Community Short Label Black Twitter btwitter Allies/Academics allies BLM Activists activists Progressives progressives Conservatives conserv Black Celebrities bcelebs Mixed Celebrities mcelebs Center/Left Politicians centleftpols Center/Left News Media centleftnews Local News localnews International Users intl
Table 4: Labeling Scheme Table translating between the long and short labels used throughout this paper.
Figure 7: Viable thresholds A knee plot showing the number of matched communities at Jaccard similarity thresholds between 0 and .3 in increments of .0125.

### v2

@@ Results
Table 2: Semantic changes: Examples of semantic changes with the tweets in which they appear, paraphrased from the original.
@@
Our approach offers insight into the changing meaning of conceptual keywords (Williams 1985) related to political activism and the BLM movement, as well as about the specific communities that helped to spread those changed meanings. Here we discuss several sets of semantic changes identified by our model and the significance of the overall network we construct. Taken together, these results affirm the validity of our model and its underlying community clustering approach, as well as the crucial work done by BLM activists to shape conversations about police violence, antiblack racism, and possible solutions to both. They also affirm the importance of Black celebrities in communicating the ideas of BLM to the general public, as well as the presence of conservative critics who push back against activist discourse — and, at times, co-opt its language. Below we discuss the findings from each phase of our methodological pipeline, building towards a conclusion about the decade-long lead-up to the current political climate in which the words of social justice and antiracism themselves have become the subject of legislative attention.
@@
The first part of our model identifies the words that demonstrated statistically-significant semantic changes (Table 2). Beginning with general terms related to activism, we find that the term activists undergoes a shift in meaning between 2014-2015, following the murders of Freddie Gray and Sandra Bland, when it has no particular connection to the work of BLM — its near neighbors are terms like “detained,” “targeted,” and “scuffle” — to the summer of 2020, by which point the term is very clearly associated with BLM activists in particular, as indicated by near neighbors including “blackled,” “movements,” “bipoc,” and “advocates.” The term action, similarly, enters the dataset with no specific connotations of activism (its near neighbors are related to TV news channels) but, by the summer of 2020, has acquired specific valences of political change, as evidenced by the near neighbors of “meaningful,” “tangible,” “steps,” and “stand.” In a similar manner, the word active also acquires connotations specific to allyship, as indicated by near neighbors of “learning,” “listening,” and “unlearning,” among others. (The terms allyship and learning undergo similar changes, acquiring each other as near neighbors along similar timeframes).
Terms like “listening” and “learning” point to more specific ways that BLM activists shaped the broader conversation about how to engage in activism around racial justice as well as how to contribute as an ally. But additional terms index specific issues and debates. For example, the term abolish begins with associations with the death penalty and, in 2020, acquires an expanded meaning that encompasses the abolition of the police — a key component of the BLM platform. The term abolishing, similarly, acquires the near neighbors of “defunding,” “reallocating,” “reforming,” and “demilitarization” as the summer of 2020 approaches. We also find theory, as in “critical race theory,” coalescing as a keyword during this time, acquiring associations with both “crt” and “ideology.” Taken together, these word changes point to a clear coalescing around the discourse of racial justice that we have come to associate with the movement.
Figure 5: Linguistic leadership: The linguistic leadership of key communities including (a) activists, (b) Black celebrities, (c) conservatives, and (d) progressives with their contributions highlighted. The height of each bar is proportional to the overall number of words a community leads (left) or follows (right). The height of the connecting band is proportional to the number of words the community on the left leads the one on right.
@@
Unexpected insights begin to emerge when looking at the leadership network overall. To be sure, we observe the influence of the BLM activists (Figure 5a), who claim an outsized number of word changes in comparison to most other communities. The progressive community, which sits slightly closer to the center than the activist community, can count a significant number of word changes as well (Figure 5d), and it would make sense that the pattern of the progressives would track that of the BLM activists but less strongly. What is perhaps unexpected is that the conservative community — moreso than the center/left news media — is the most substantial follower of both the activist and progressive communities. While we might expect a more gradual path from the left to the center, and then to the right, our method shows how the conservative community engaged directly with the activist and progressive communities from the start.
@@
Another unexpected but not surprising result was the strong role of the community of Black celebrities in communicating the ideas of the BLM movement to the general public (Figure 5b). Indeed, Black celebrities claim 24.73% of all of word changes, and are followed by activists, then center/left news media, and conservatives in high proportion. This finding suggests that the direct association with the BLM movement may mark its activists for pushback — and, at times, outright aggression — by the conservative community, the celebrity status of those in the Black celebrity community may protect them from some of the conservative vitriol, and as a result, allow them to communicate the ideas of the activists to a broader public. We see this in their leadership of terms like activism, which transforms from the specific context of social movements to a word that itself indexes a role of social importance, as indicated by the near neighbors of “importance,” “career,” “societal,” and “impact.” We also see this in the terms that one would seem to associate with the BLM movement itself, such as listen. The connotations of learning and unlearning, as well as amplifying the voices of those with direct experience-as indicated by the near neighbors of “learn,” “unlearn,” “amplify,” and “explain,” are communicated to the center/left news media commmunity from the Black celebrities. In addition, we find this community influencing the community of progressives, passing along the changed meaning of theory, discussed above, as well as the more basic concept of equity. Interestingly, the conservative community at times learns directly from the Black celebrity community. In particular, the term nonracist is introduced by this community, and is adopted directly from the conservative community from there.
@@
A final observation has to do with the role of the conservative community in introducing new word meanings-and not just adopting them (Figure 5c). While few of these words are conceptual keywords, their overall number is substantial. This finding confirms the consistent and persistent engagement of these ideas by the conservative community — not only in the wake of the summer of 2020, but from the movement’s start. That these words are not conceptual keywords at once confirms prior findings about how conservatives — as well as white people as a group — demonstrate a lead-lag relationship with BLM activists (Le-Khac, Antoniak, and So 2023) and shows how they engaged at the level of discourse: through the wide array of words that make up a language.
@@
As far as the prominence of the conservative community in our study, we see significant implications for our understanding of the emergence of the “anti-woke” and “anti-DEI” agenda that, at the time of this writing, has entered legislation in nine US states (Ray and Gibbons 2021) and appears poised to wholly reshape government agencies and funding at the federal level. Whereas the common narrative is that this movement emerged in the aftermath of the summer of 2020, and has since been fueled by the Republican party, our findings show that there has always been direct conservative engagement with the discourse of Black Lives Matter. Moving forward, this finding may reshape the narrative of “conservative backlash” and, instead, point to a story of continual pushback and even outright aggression that met the movement from its start.
Finally, we would like to highlight a key absence in this project, which is the missing data that might otherwise document the crucial role of the young Black activists who fueled the movement from its start. These regular young people were essential in the movement’s early phases, as they were in organizing the on-the-ground protests that took place in every city where a police killing took place. Qualitative researchers have documented their crucial role (Jackson, Bailey, and Welles 2020; Clark 2024), as have news reports (Xue 2020), yet their status as regular young people has meant that they lack the protections granted to celebrities or even to adults. As activists, they have been subjected to harassment and intimidation, and in some cases they have even been killed by the police themselves (Press 2019; del Rio 2020). As early as mid-2015, when Freelon et al released their initial findings on the network structure of #BLM, they found that the young Black activists who had been present in their data in 2014 had disappeared eighteen months later (Freelon, McIlwain, and Clark 2016). Using the usernames of their young Black activist community as seeds, we found similarly that these users had largely disappeared-save for the few who themselves became famous, and joined the Black celebrity community instead.
In her influential theorization of missing data sets, the artist and educator Mimi Onuoha explains how the gaps in datasets provide “cultural and colloquial hints of what is deemed important,” and further observes how “this lack of data typically correlates with issues affecting those who are most vulnerable in that context” (Onuoha 2018). In our #BlackLivesMatter dataset and the findings that it has enabled, these young Black activists constitute this precise form of missing data. Thus as we celebrate our findings for how they affirm the role of the Black Lives Matter movement in shaping a larger public conversation, we must also hold space for what our study cannot confirm, which is the equally crucial role played by the activists whose words are missing from our dataset. Even if their words are unrecorded, they are not unremembered.
@@ Conclusion
In this project we describe a method of modeling semantic leadership across a set of communities associated with the Black Lives Matter movement, which has been informed by qualitative research on the structure of social media and of Black Twitter in particular. We describe our bespoke approaches to time-binning, community clustering, and connecting communities over time, as well as our adaptation of state-of-the-art approaches to semantic change detection and semantic leadership induction. We find substantial evidence of the leadership role of BLM activists and progressives, as well as Black celebrities. We also find evidence of the sustained engagement of the conservative community with this discourse, suggesting an alternative explanation for how we have arrived at the present political moment, in which “anti-woke” and “anti-DEI” policy is being enacted nation-wide. Contrary to the dominant narrative of conservative backlash against a movement that became too radical for the center to support, we find that the conservative community had been engaging directly with the BLM movement from its very start. The conservative community is the largest follower of word changes across the entire timespan of our study, and introduces many word changes as well, albeit few original conceptual keywords.
From our present vantage-point, in Spring 2025, we can now see the result of this sustained conservative engagement: a distortion of the meaning of the words that Black Lives Matter worked so hard to bring to public attention, ultimately weaponizing the most powerful of these words to lead to policies that dismantle, rather than more fully realize, the movement’s original goals. With this in mind, just as we recognize the role of the young Black activists who are missing from our data, we also recognize and hold space for the valuable work of all those involved in the Black Lives Matter movement. While the immediate gains of their work are currently being dismantled, the ideas behind their words remain a powerful force. We submit this paper as evidence of what words can accomplish when coupled with a commitment to action and an unwavering belief in the importance of equality and justice for all.
Acknowledgments
Thank you to Tanvi Sharma for her design work on the final visualizations. This project would not have been possible without the data collection work of Alex Fan. It also greatly benefited from discussions with André Brock and Jacob Eisenstein. This project has been supported by a grant from the Mellon Foundation (G-2211-14240) and by Emory’s Department of Quantitative Theory & Methods.
@@ Acknowledgments
Maxima Additional Candidate Maxima
Sep 02, 2014 n/a
Dec 25, 2014 Dec 18, 2014
Aug 31, 2015 Mar 14, 2015; Apr 10, 2015; May 04, 2015
Jul 08, 2016 n/a
May 27, 2020 n/a
Aug 07, 2020 Sep 05, 2020; Sep 25, 2020
Table 3: Maxima: Additional candidate maxima that were merged into the final time bin segments.
Community Short Label
Black Twitter btwitter
Allies/Academics allies
BLM Activists activists
Progressives progressives
Conservatives conserv
Black Celebrities bcelebs
Mixed Celebrities mcelebs
Center/Left Politicians centleftpols
Center/Left News Media centleftnews
Local News localnews
International Users intl
Table 4: Labeling Scheme: Table translating between the long and short labels used throughout this paper.
Figure 6: Top Community Members: The most central users of each subcommunity cluster, with color indicating final community grouping. Users with fewer than 3000 followers (as of September 2024) or with inactive accounts are not shown. Note that ranking by in degree leads to certain users (e.g. @realdonaldtrump) being included in communities in which they engage adversarially.
Figure 7: Viable thresholds: A knee plot showing the number of matched communities at Jaccard similarity thresholds between 0 and .3 in increments of .0125.